### v3

# The double gamma function and Vladimir Alekseevsky

The double gamma function and Vladimir Alekseevsky
Yury A. Neretin
This paper is about a forgotten function and a forgotten mathematician. The double gamma function is now an important special function, which appears for different reasons in many branches of mathematics and in mathematical physics, as it satisfies a large and amusing collection of identities parallel to the classical gamma function. It was discovered and investigated in detail by Vladimir Alekseevsky in 1888-89. We outline his starting point here: considering the Weierstrass product for the entire periodic function $\sin\pi x$ and taking half of the factors corresponding to non-positive roots of the sine, we obtain the function $1/\Gamma(x)$. Considering the Weierstrass product for the doubly quasiperiodic Jacobi theta function $\vartheta_{1}$ and taking the half of factors we come to the $q$-gamma function (which appears, with a slight change in notation, in Eduard Heine’s work). Considering the quarter of the factors (corresponding to the positive quadrant in the lattice $n_{1}\omega+n_{2}\omega_{2}$ of periods), Alekseevsky arrived at his double gamma function. The work in this direction was continued by Ernest Barnes, Jean Beaupin, Godfrey Hardy, and Vladimir Steklov in 1899–1907. After 1907, publications on this topic stopped, and the double gamma appeared again 70 years later due to Takuro Shintani. In this paper we discuss Alekseevsky’s seminal work and its genesis, the history of the double gamma, and Alekseevsky’s biography (1858-1916).
AMS classification: 01A55, 01A70, 33B15, 33E99, 30D05.
Key words. Mathematics of XIX century, mathematics of XX century, special functions, gamma function, theta function, complex analysis, functional equations, Weierstrass factorization theorem, universities, biography.
4000000000000000000000000.Content.
1. Introduction. A forgotten function and a forgotten mathematician
2. Biography of Alekseevsky
3. Alekseevsky’s work, 1888-89
4. On the history of the double gamma function
5. Alekseevsky: historiography, lost work, and publications
6. Conslusion.
4.References
1. Introduction. A forgotten function and a forgotten mathematician
In 1996, Vladimir Fock spoke to me about the work [Fad1995], [Fock1997], [Kash1998], [FoCh1999] being done on the quantization of Teichmüller spaces; this construction was essentially based on a couple amusing special functions called ’Barnes double sine’ and ’Barnes double gamma’. I read the original papers of Ernest Barnes [Bar1899][Bar1904] and observed that he referred to an earlier work of Alexeiewsky. This name was unknown to me, and my immediate attempt to find an information about this person and his works was unsuccessful.
Many years later, collecting material for a book about Nikolay Luzin, I opened Krulikovsky’s ’History of the development of mathematics in Tomsk[Kru1967]. Here I was again met with the name Aleksevsky. I found his main work and was very surprised reading it. I then started a historical excavation, which I now present to the reader.
The double gamma function is now a well recognized special function used in different branches of mathematics and in mathematical physics. It is an entire function $G(x;\alpha)$ of a complex variable $x$ and depending on a complex parameter $\alpha$, which satisfies the following functional relations:
(1.1) $G(x+1;\alpha)$ $=\Gamma\Bigl(\frac{x}{\alpha}\Bigr)\,G(x;\alpha);$
(1.2) $G(x+\alpha;\alpha)$ $=(2\pi)^{\frac{\alpha-1}{2}}\alpha^{-\frac{2x-1}{2}}\,\Gamma(x)\,G(x;\alpha),$
and is normalized by
$G(0;\alpha)=1.$
This function satisfies a big collection of properties parallel to the classical gamma function.
A special case of $G(x,\alpha)$ corresponding to $\alpha=1$ was discovered by Hermann Kinkelin in [Kin1860]; notice that in this case the two equations (1.1), (1.2) coincide. His remarkable work did not attract interest and was forgotten for a long time. The double gamma for general $\alpha$ was introduced by Vladimir Alekseevsky [^1] Russian: Алексеевский; old Russian orthography: Алексевскiй. I use the most natural English spelling of the name, which is used by his modern homonyms. German and French transliterations used in 1884-1907 were Alexeiewsky, Alexejewsky, Alexéievsky. in a large memoir [Ale1888][Ale1889]. This work was continued by Ernest Barnes [Bar1899], [Bar1900.1], [Bar1900.2], [Bar1901], [Bar1903], [Bar1904], Jean Beaupain [Bea1903], [Bea1904], [Bea1907], Godfrey Hardy [Har1905.1], [Har1905.2], and Vladimir Steklov [Ste1904]. The chain of publications on the double gamma and the multiple gamma was interrupted in 1907, and after which they were more or less forgotten for 70 years.
Alekseevsky (1858–1916), the discover of the double gamma, was also forgotten. He was mentioned in a few Russian papers on the histories of mathematics and education, [Gus1950], [Gus1952], [Kru1967], [Bakh1967], [Kru1967.1], [Yush1968], [Gran2003], [Bel2000] (mainly, on the local history of mathematics in Kharkov and Tomsk). He studied and worked at Kharkov University (1877-1905) and Tomsk Technological Institute (1905-1916). For the Kharkov period, we mainly have records of unsuccessful attempts to advance his career. In Tomsk he became a full professor of mechanics, and in 1907-1911 he was the director (rector) of the Institute.
We briefly discuss his biography in Sect. 2. But our main topics are his impressive central work and the difficulty of mathematical development.
In Sect. 3, we discuss Alekseevsky’s main work [Ale1888][Ale1889] and its genesis. It was a non-obvious response to Eduard Heine’s investigations of $q$-hypergeometric functions (though, more precisely, Alekseevsky’s starting point was Paul Appell’s work [App1881] on extending the $q$-gamma).
In Sect. 4, we discuss the history of the double gamma. We start with a brief exposition of Kinkelin’s work. It was basically surpassed in [Ale1888][Ale1889], but it also contains important results, which were not repeated in Alekseevsky’s memoir (the link between Kinkelin and Alekseevsky was later observed by Beaupin). We also try to understand the origin of Kinkelin’s work. It seems to have been an investigation in a ’speleological’ style, which is relatively common in theory of special functions.
The forgetting of important research [^2] Opening the database MathSciNet, which is quite selective, I see 233 references to [Bar1899] starting 2002. At the very least, this is unusual for a paper published a hundred years before. Google Scholar shows 512 references to [Bar1904]. is in itself an interesting historical phenomenon. In this case, we see impressive articles clearly written in a common mathematical language; they were available to readers of the XX century as well as readers of 1890-00s. We are trying to understand this phenomenon.
Sect. 5 contains three additions. Firstly, we briefly discuss historical sources on Alekseevsky and historiography.
Secondly, starting with [Ale1892], [Ale1894] Alekseevsky claimed many times over his researches could be extended to a wider class of functions depending on several parameters. We analyze these claims — clearly, he kept in mind the multiple gamma functions; definitely he did not publish theorems on this topic. However, these claims were known and made water run in circles. Apparently, [Bar1904] was a response to a declaration in [Ale1894].
To be complete, I present a list of his textbooks on calculus and mechanics published by universities (all research publications are listed in Bibliography, [Ale1884.1][Ale1904]; I think that this list is complete).
2. Biography of Alekseevsky
2.1. Early biography. Vladimir Petrovich Alekseevsky was born April 19 (Julian 07), 1858 in Novgorod governorate [^3] Russian: ‘губерния’ (‘guberniya’). and baptized in the village Peretënka located near Okulovka station on the railway line Moscow–St.Petersburg. His father, Petr Osipovich Alekseevsky, was a ‘praporshchik’ $\simeq$ warrant officer or sergeant major] of the Corps of transport engineers [Корпус инженеров путей сообщения], [Sin1917], [Gran2003].
In 1877, Alekseevsky graduated from the Nizhny Novgorod gymnasium. It was a good school, with many of its pupils going on to become well known, in particular, mathematicians Aleksandr Lyapunov and Ivan Privalov, physicist Grigory Landsberg [^4] Landsberg in particular could have won the Nobel in 1930 (combinational scattering of light), but was not nominated for it. He also edited the famous high-level ¡¡Elementary physics textbook¿¿, which was used by several generations of Russian natural scientists, engineers, and teachers., zoologist and ecologist Aleksandr Formozov, philosophers Semyon Frank and Vasily Rozanov, and chemist Alexey Favorsky.
2.2. Kharkov University, 1877-1905. After graduating from the gymnasium, Alekseevsky entered the Natural Science Department of the Faculty of Physics and Mathematics of Kharkov University. Two years later, he transferred to the Mathematical Department. In 1881, ‘for family reasons’, he left the university [^5] The main obstacle to his research activity was his difficult financial situation’ as was written in [Bakh1967]; this is also clear from Sintsov [Sin1917]). However, Sintsov in the obituary does not provide any details (and they are absent in other sources).. In 1884, he passed examinations without attending lectures and graduated from the university. At that moment he had three published papers [Ale1884.1]-[Ale1884.3] on the integrability of certain ordinary differential equations. They were considered as a graduate (‘candidate’) work [^6] There is no information about his scientific adviser(s). In [Ale1884.1]-[Ale1884.3], he mentioned the names of Vasily Imshenetsky (1832-1892), who worked at the Kharkov University 1877-1881 (and moved to St.-Petersburg afterward), and Aleksey Letnikov (1837—1888), who worked in Moscow..
Since 1881, Alekseevsky taught mathematics, mechanics, and physics in Kharkov in a real school [^7] A high school oriented to natural sciences, engineering, commerce, mathematics. (until 1891) and in the Second women’s gymnasium (until 1887). In 1887, he received a scholarship to prepare for a professorship (analogous to becoming a postgraduate student in the present).
Alekseevsky’s main work On functions similar to the gamma function was published in Communications of Kharkov Mathematical Society [^8] A well-known journal between 1879-1918. Russian: Сообщения и протоколы заседаний Математического общества при Императорском Харьковском университете; old Russian orthography: Сообщенiя и протоколы засданiй Математическаго общества при Императорскомъ Харьковскомъ университет; French: Communications et procus-verbaux de la société mathématique de Kharkow; German: Sammlung der Mitteilungen und Protokolle der Mathematischen Gesellschaft in Charkow. [Ale1888][Ale1889] (in fact, it is one paper split in the middle of a sentence). He defined the double gamma function $G(x;\alpha)$ satisfying two functional equations and derived an unexpectedly large collection of its properties. We discuss this impressive work and its genesis in Sect. 3.
In 1892, this work was presented as a dissertation [^9] A ‘magister dissertation’ in the terminology of that time. In modern terminology, it apparently corresponds to a PhD. [Ale1892]. After the defence (21.02.1893), he became a privatdozent in Kharkov University.
In March 1894 – November 1895 Alekseevsky visited Leipzig, Göttingen, and Paris. At that time, he published a short paper [Ale1894] in German in Leipz. Ber.. Apparently, his work became known due to this publication.
After coming back, he decided to leave the university and teached at a gymnasium in town Starobelsk, in Kharkov governorate since ’university …could do nothing to support the scientist it highly estimated’ (a quotation of a rapport of Vladimir Steklov in [Sin1917], who worked at Kharkov university in that time). Alekseevsky returned to the university the next year. Starting 1898, he also worked at Kharkov Technological Institute.
Four pages of his obituary [Sin1917] (and a big part of [StSi1904], see, also, the correspondence of Steklov and Aleksandr Lyapunov in [Ste1991]) contain a description of the efforts of several colleagues, through 1896-1904, to transfer Alekseevsky to the position of extraordinary professor. These included Aleksandr Lyapunov, Vladimir Steklov, Matvey Tikhomandritsky (specialist in elliptic functions), Dmitry Grave, Konstantin Andreev (later a professor and dean in Moscow), Ludwig Struve (astronomer [^10] Grandson of the famous astronomer Friedrich-Vasily Struve.), and Ivan Osipov (chemist). The nature of obstacles is not clear from the published literature. Steklov (letter to Lyapunov, 14.11.1903) characterized this as an ’absurd story’.
From [Sin1917]: Steklov [1902] explains, why V. P.’s doctoral dissertation $\simeq$ habilitation] was delayed – ’he could have compiled from the scientific results a fully satisfactory dissertation’ and expresses his conviction that ‘if V. P. did not take care to acquire formal rights, having at his disposal all the scientific materials, this circumstance can hardly diminish the scientific merits of V. P. Alekseevsky …But we must say even more. By now, V. P. Alekseevsky has prepared for printing and is already printing an extensive course of differential and integral calculus.…V. P. Alekseevsky’s course is not a template course adapted to the usual programs of higher educational institutions. It is an independent academic work, where, along with pedagogical requirements, it combines a rigorous and modern approach to the study of the subject’ (the book in question was not published).
After fifth attempt, 12.01.1904 Alekseevsky got the position of acting extraordinary professor in Kharkov. In the next year, he moved to Tomsk for a full professorship.
It seems doubtless (see Subsect. 3 and Subsect. 5) that Aleksevsky had a continuation of his work devoted to the multiple gamma functions and the multiple zeta functions. However, for unknown reasons he has not published these investigations. It seems also that this was one of the sources of his career difficulties (see footnote 23 below).
2.3. Tomsk Technological Institute, 1905–1916. The institute was founded in 1900 by chemist-technologist Efim Zubashev (see [Gal2004], [Gal2009]) and quickly became a center of education and science in Siberia. For instance, among its professors in the beginning of the XX century were Theodor Molien — one of the founders of representation theory of finite groups (together with George Frobenius and William Burnside) and theory of associative algebras [^11] See Bourbaki [Bou1960], Section ‘Noncommutative algebra’, van der Waerden [Wae1985], and selected works of Molien in [Mol1985]. — and the famous geologist Vladimir Obruchev.
According [Gran2003], Zubashev invited Alekseevsky to Tomsk for first time in 1900 (the Institute’s the first year of work). In April 1905, director Zubashev offered Alekseevsky the position of acting ordinary professor (full professor) of theoretical mechanics in Tomsk. Alekseevsky agreed and, starting 28.11.1905, worked in Tomsk. He left pure mathematics, and I could not find any his scientific publications by him of that time.
Granina’s paper [Gran2003] contains a lot of data on the educational and administrative activity of Alekseevsky in Tomsk. We mention only a few formal details.
Students of Tomsk Technological Institute were involved in the tumultuous events of the First Russian Revolution of 1905-07. Zubashev (who worked under pressure from both progressists and conservators) was accused of being loyal to the striking students and was exiled from Tomsk. Later, he returned to the director position but had to leave it ‘for health reasons’ (November 1907). The institute council elected Alekseevsky as the new director (rector), and his administrative activity was considered in local annals as quite successful and fruitful.
In 1911, there was a wave of student and professor protests in Russia against the education reform initiated by minister Lev Kasso. The Institute was involved in these protests. Several professors were dismissed for oppositionism [^12] The list of fired professors includes Molien and Obruchev. (8.06.1911). Aleksevsky and the institute council tried to reinstate them. As a result, by an imperial edict, Alekseevsky was ousted from the director position. However, he continued his full professorship and some administrative activity.
In 1916 he stubbed his toe, didn’t treat it, and died of gangrene on 13 May.
3. Alekseevsky’s work, 1888-89
Alekseevsky’s work On functions similar to the gamma function was published in Communications of Kharkov Mathematical Society in 1888-1889 in two papers [Ale1888], [Ale1889].
3.1. The function $G(x)$. The Alekseevsky’s paper begins with the words: The first problem, whose solution is the topic of the present investigation, is to examine the properties of a function $G(x)$ satisfying the equation:
(3.1) $G(x+1)=\Gamma(x)\,G(x)$
under the condition
(3.2) $G(1)=1.$
There are no references, no names, no motivations, no acknowledgements here. It is hard to believe that the author simply decided to attack such an exotic problem and successfully solved it. In reality, the presentation of the paper was backwards to the real way of thinking; this is discussed below in Subsect. 3.
Taking the logarithm of both sides of (3.1), we come to the difference equation
$\ln G(x+1)-\ln G(x)=\ln\Gamma(x).$
According to the Malmsten formula, see for example [Erd1953], formula 1.9(1),
$\ln\Gamma(x)=\int_{0}^{\infty}\frac{e^{-u}\,du}{u}\Bigl\{(x-1)-\frac{1-e^{-(x-1)u}}{1-e^{-u}}\Bigr\}.$
Denote the integrand by $L_{1}(x,u)$. Alekseevsky (§1) writes the simplest function $L_{2}(x,u)$ such that
(3.3) $L_{2}(x+1,u)-L_{2}(x,u)=L_{1}(x,u)$
and defines $\ln G(x)$ as $\int_{0}^{\infty}L_{2}(x,u)\,du$,
(3.4) $\ln G(x):=\int_{0}^{\infty}\frac{e^{-u}\,du}{u}\Bigl\{\frac{(x-1)(x-2)}{2}-\frac{x-1}{1-e^{-u}}+\frac{1-e^{-(x-1)u}}{(1-e^{-u})^{2}}\Bigr\}.$
Clearly, $G(x)$ satisfies (3.1) and
(3.5) $G(n+1)=\frac{1^{1}2^{2}\dots n^{n}}{(n!)^{n}}=1^{n}\cdot 2^{n-1}\cdot(n-1)^{1}.$
Clearly, a solution of the functional equation (3.2) is determined up to a periodic factor. Further results show that it is a ‘correct’ solution in an informal sense.
In §5, using the integral representation (3.4), he gets a decomposition of $G(x)$ into a Weierstrass product:
(3.6) $G(1+x)=(2\pi)^{x/2}e^{-\frac{x(x+1)}{2}-\frac{\gamma}{2}x^{2}}\prod_{n=1}^{\infty}\Bigl\{\Bigl(1+\frac{x}{n}\Bigr)^{n}e^{-x+x^{2}/(2n)}\Bigr\},$
where
$\gamma:=\lim_{n\to\infty}\Bigl(1+\frac{1}{2}+\dots\frac{1}{n}-\ln n\Bigr)$
is the Euler constant (cf. the similar expression for $\Gamma(x)$ in [Erd1953], 1.1 (3)). Next, he notices that this formula gives an expression for $G(x)$ on the whole complex plane.
In §2, Alekseevsky gets a counterpart of the Euler definition of the gamma function, cf. [Erd1953], 1.1 (2),
(3.7) $G(x)=\lim_{n\to\infty}\Bigl\{(n+1)^{\frac{(x-1)(x-2)}{2}}n!^{x-1}\prod_{k=0}^{n-1}\frac{\Gamma(k+1)}{\Gamma(k+x)}\Bigr\}.$
The counterpart of the Legendre duplication formula (cf. [Erd1953], 1.2(15)) is (§9):
$G(2x)=\frac{1}{G(1/2)^{2}}2^{(x-1)(2x-1)}\,\pi^{-x}\,\Gamma(x)\,G^{2}(x)\,G^{2}(x+\tfrac{1}{2}).$
After complicated calculations with functions $G(x;\alpha)$ in §21 (see below Subsect. 3) he gets the multiplication formula (cf. [Erd1953], 1.2(11)) in the form
(3.8) $G(nx)=\frac{n^{\frac{(nx-1)^{2}}{2}}}{(2\pi)^{\frac{(n-1)(nx-1)}{2}}}\,\,\prod_{j=1}^{n-1}\frac{\Gamma\bigl(x+\frac{j-1}{n}\bigr)^{n-j}}{\Gamma\bigl(\frac{j}{n}\bigr)^{n-j}}\,\,\prod_{j=0}^{n-1}\frac{G\bigl(x+\frac{j}{n}\bigr)^{n}}{G\bigl(\frac{1+j}{n}\bigr)^{n}}.$
The multiplication formula admits a simpler form obtained by Kinkelin [Kin1860], see formula (4.2) below.
In §11, Alekseevsky establishes the following counterpart of the ’Stirling’ asymptotic expansion (cf. [Erd1953], 1.18(1)):
$\ln G(x+1)\sim\ln\frac{\pi^{1/6}G(\frac{1}{2})^{\frac{2}{3}}}{2^{\frac{1}{36}}}+\frac{x}{2}\ln 2\pi+\Bigl(\frac{x^{2}}{2}-\frac{1}{12}\Bigr)\ln x-\frac{3}{4}x^{2}+\\ +\sum_{n=1}^{\infty}\frac{(-1)^{n}B_{2n+2}}{4n(n+1)x^{2n}},\qquad x\to+\infty.$
where $B_{2n}$ are the Bernoulli numbers, $\frac{x}{e^{x}-1}=\sum_{n}\frac{B_{n}}{n!}x^{n}$.
3.2. Integrals and products. Next, he evaluates some integrals in terms of the $G$-function,
(3.9) $\int_{0}^{a}\ln\Gamma(x)\,dx=-\ln G(a)+(a-1)\ln\Gamma(a)-\frac{a(a-1)}{2}+\frac{a}{2}\ln 2\pi,$
this identity (§10) can be regarded as an integral representation of $\ln G(x)$.
Keeping in mind the reflection formula for the gamma function, we come (§14) to
(3.10) $\int_{0}^{x}\ln\sin\pi x\,dx=x\ln\frac{\sin\pi x}{2\pi}+\ln\frac{G(1+x)}{G(1-x)}.$
This allows us to get some other integrals as
(3.11) $\int_{0}^{x}\pi x\cot\pi x\,dx=x\ln 2\pi+\ln\frac{G(1-x)}{G(1+x)}.$
We also mention the following formula (§5),
$\prod_{k=0}^{n-1}\frac{G(a-e^{2\pi ik/n}x)}{G(a)}=\prod_{m=0}^{\infty}\Bigl(1-\frac{x^{n}}{(a+m)^{n}}\Bigr)^{m+1}.$
3.3. The multiple gamma functions. Denote
$G_{1}(x):=\Gamma(x),\qquad G_{2}(x):=G(x).$
In §16, he introduces functions $G_{n}(x)$ satisfying the relations
(3.12) $G_{n}(x+1):=G_{n-1}(x)\,G_{n}(x),\qquad G_{n}(1)=1.$
Namely,
(3.13) $\ln G_{n}(x)=\int_{0}^{\infty}\frac{e^{-u}du}{u}P_{n}(x),$
where
(3.14) $P_{n}(x+1)=\sum_{i=0}^{n}(-1)^{i}\frac{[x]_{n-i}}{(n-i)!}\frac{1}{(1-e^{-u})^{i}}+(-1)^{n+1}\frac{e^{-xu}}{(1-e^{-u})^{n}},$
and
$x]_{k}:=x(x-1)\dots(x-k+1).$
The expressions $P_{n}(x)$ satisfy the difference equation
$P_{n+1}(x+1)-P_{n+1}(x)=P_{n}(x),$
which implies the desired property.
Alekseevsky does not develop a further theory of such functions, but in §17-19 he discusses some generalities of function theory related to the Weierstrass factorization theorem. Namely, he considers arbitrary products of the form
$f_{0}(z)=\prod_{k=0}^{\infty}\Bigl(1+\frac{z}{a_{k}}\Bigr),\qquad\text{where $\mathop{\mathrm{Re}}\nolimits a_{k}>0$, $\sum|a_{k}|^{-1}<\infty$}.$
He represents $\ln f_{0}(z)$ as a Laplace transform, and, applying the same trick with functions $P_{n}(z)$, constructs a sequence of functions $f_{1}(z)$, $f_{2}(z)$, …such that
$f_{n}(z+1)=f_{n}(z)f_{n+1}(z).$
He writes explicit representations of functions $f_{n}$ as products of functions $G_{m}$ with given $m\leqslant n$.
3.4. Double gamma functions with two periods. In §20, Alekseevsky writes: Functions similar to $\Gamma(x)$ considered above are special cases of more general functions. Since they are related to theta-functions, we derive some properties of one such function.
Fix a parameter $\alpha$ (which initially is positive real, but later it becomes complex). Alekseevsky considers a function [^13] He uses the notation $H(x)$ or $H(x,\alpha)$. $G(x;\alpha)$ such that
$G(x+1;\alpha)=\Gamma\Bigl(\frac{x}{\alpha}\Bigr)\,G(x;\alpha),\qquad G(1;\alpha)=1.$
This function is defined by
(3.15) $\ln G(x;\alpha)=\\ =\int_{0}^{\infty}\frac{e^{-\alpha u}du}{1-e^{-u}}\Bigl\{\frac{(x-1)(x-2\alpha)}{2\alpha}-\frac{x-1}{1-e^{-\alpha u}}+\frac{e^{-(1-\alpha)u}-e^{-(x-\alpha)u}}{(1-e^{-u})(1-e^{-\alpha u})}\Bigr\}$
It turns out (this is non-obvious) that $G(x;\alpha)$ satisfies the second functional equation
$G(x+\alpha;\alpha)=(2\pi)^{\frac{\alpha-1}{2}}\alpha^{-\frac{2x-1}{2}}\Gamma(x)\,G(x;\alpha).$
Alekseevsky finds the main properties (cf. [AlKu2023]) of this function, in particular,
$G\Bigl(x;\frac{1}{\alpha}\Bigr)=\frac{G(\alpha x;\alpha)}{G(\alpha,\alpha)^{x}}\,\alpha^{\frac{(x-1)(\alpha x-2)}{2}},$
see §21. In fact (§20),
$G(\alpha,\alpha)=\alpha^{-\frac{1}{2}}(2\pi)^{\frac{\alpha-1}{2}}.$
Next (§21),
$\frac{G(x;\alpha)}{G\bigl(\frac{x}{\alpha+1};\frac{1}{\alpha+1}\bigr)\,G\bigl(\frac{x}{\alpha+1};\frac{\alpha}{\alpha+1}\bigr)}=G(\alpha,\alpha)^{\frac{x}{\alpha+1}}\,\alpha^{\frac{-x(x-\alpha-1)}{2(\alpha+1)}}(\alpha+1)^{\frac{x-\alpha-1}{\alpha+1}}\Gamma\Bigl(\frac{x}{\alpha+1}\Bigr).$
If $m$, $n$ are positive integers, then $G\bigl(x,\frac{m}{n}\alpha\bigr)$ can be expressed in terms of $G(x;\alpha)$. Namely,
$G\Bigl(x,\frac{m}{n}\alpha\Bigr)=\frac{n^{\frac{(x-1)(nx-m\alpha)}{2m\alpha}}}{(2\pi)^{\frac{(n-1)(x-1)}{2}}}\,\,\prod_{k=0}^{n-1}\prod_{j=0}^{m-1}\,\frac{G\bigl(\frac{nx+nj+km\alpha}{mn};\alpha\bigr)}{G\bigl(\frac{n+nj+km\alpha}{mn};\alpha\bigr)},$
see §21. This formula has numerous interesting special cases, for instance, we can set $m=1$, or $n=1$, or $\alpha=1$, or $\alpha=1/n$, or $\alpha=n$, or $m=n$. The multiplication formula (3.8) also is a non-obvious corollary of the last formula.
In §22, he obtains two limit expressions for $G(x;\alpha)$ extending (3.7),
$G(x;\alpha)$ $=\lim_{n\to\infty}\Bigl(1+\frac{n}{\alpha}\Bigr)^{\frac{(x-1)(x-2\alpha)}{2\alpha}}\Gamma\Bigl(1+\frac{n}{\alpha}\Bigr)^{x-1}\,\,\prod_{k=0}^{n-1}\frac{\Gamma\bigl(\frac{1+k}{\alpha}\bigr)}{\Gamma\bigl(\frac{x+k}{\alpha}\bigr)}$
$=\lim_{n\to\infty}\alpha^{n(x-1)}\Bigl(\frac{1}{\alpha}+n\Bigr)^{\frac{(x-1)(x-2\alpha)}{2\alpha}}\Gamma\Bigl(\frac{1}{\alpha}+n\Bigr)^{x-1}\,\,\prod_{k=0}^{n-1}\frac{\Gamma(1+k\alpha)}{\Gamma(x+k\alpha)}.$
Also, he obtains a Weierstrass type decomposition (§22),
(3.16) $G(x;\alpha)=e^{ax+bx^{2}}\frac{x}{\alpha}\times\\ \times\prod_{m\geqslant 0,n\geqslant 0;(m,n)\neq(0,0)}\Bigl(1+\frac{x}{m+n\alpha}\Bigr)\,\exp\Bigl\{-\frac{x}{m+n\alpha}+\frac{x^{2}}{(m+n\alpha)^{2}}\Bigr\},$
where the constants $a$, $b$ are defined by
(3.17) $a:$ $=\Bigl(\frac{\partial}{\partial x}\ln G(x;\alpha)+\frac{1}{\alpha}\frac{d}{dx}\ln\Gamma(x)\Bigr)\biggr|_{x=1};$
(3.18) $b:$ $=\frac{1}{2}\Bigl(\frac{\partial^{2}}{\partial x^{2}}\ln G(x;\alpha)+\frac{1}{\alpha^{2}}\frac{d^{2}}{dx^{2}}\ln\Gamma(x)\Bigr)\biggr|_{x=1}.$
3.5. How was the double gamma function invented? After formula (3.16), Alekseevsky writes: Taking this product as a definition of $G(x,\omega)$ and setting
$x=\frac{z}{\omega},\qquad\alpha=\frac{\omega^{\prime}}{\omega},$
where $\omega$, $\omega^{\prime}$ are periods of elliptic functions, we get
(3.19) $G\Bigl(\frac{z}{\omega};\frac{\omega^{\prime}}{\omega}\Bigr)=e^{a\frac{z}{\omega}+b\frac{z^{2}}{\omega^{2}}}\frac{z}{\omega^{\prime}}\times\\ \times\prod\Bigl(1+\frac{z}{m\omega+n\omega^{\prime}}\Bigr)\exp\Bigl\{-\frac{z}{m\omega+n\omega^{\prime}}+\frac{z^{2}}{(m\omega+n\omega^{\prime})^{2}}\Bigr\}.$
It is clear that this function consists of a quarter of the factors entering to the $\varsigma_{3}$-function of Weierstrass [^14] The Weierstarss function, see for example [Akh1970], §13, is given by a product (3.19) over $(m,n)\in\mathbb{Z}^{2}\setminus(0,0)$. In our case (3.16), the product is given over $m\geqslant 0$, $n\geqslant 0$, $(m,n)\neq(0,0$. (or $\vartheta_{1}$ of Jacobi), herein it comes from $\varsigma_{3}$ the same way as $\frac{1}{\Gamma(z)}$ comes from $\sin\pi z$. However, in a two-periodic function, the ratio $\frac{\omega^{\prime}}{\omega}$ can be imaginary or complex but for the function $G$ this condition is not necessary151515i.e., a ratio of periods can be positive real..
The function $G$ is analogous to the function of Heine [^16] Footnote from [Ale1888]: Handbuch der Kugelfunctionen, p. 109, or Crelle J. t 34. p. 290, see [Hei1847], [Hei1878]. I do not understand these references (they are taken from [App1881]). However, the attribution of this expression to Eduard Heine is correct, see the next subsection., which is composed from a half of the factors of $\vartheta_{3}$. For a collation, we present an expression from Appell’s memoire [^17] Footnote from Alekseevsky: Mathematische Annalen. B. 19. p. 84. , see [App1881],
$O(x,\omega,\omega^{\prime})=e^{ax^{2}+bx+c}\,x\prod_{m\in\mathbb{Z},n\geqslant 0,(m,n)\neq 0}\Bigl(1-\frac{x}{n\omega_{1}-m\omega_{2}}\Bigr)e^{\frac{x}{n\omega_{1}-m\omega_{2}}+\frac{x^{2}}{(n\omega_{1}-m\omega_{2})^{2}}}$
for
$m=0,\,\pm 1,\,\pm 2,\,\dots$$\pm\infty$;
$n=0,\,+1,\,+2,\dots,+\infty,$
excluding the combination $m=0$, $n=0$.
This explains the origin of this work. Looking to the formula in Appell’s paper, Alekseevsky understood that it is possible to manipulate with Weierstrass-type product over a quadrant in the lattice $\mathbb{Z}^{2}$. However, he found that the integral representation (3.4) is the key to the whole story, and the initial definition of the work became the final theorem of his published paper (in fact, an evaluation of constants $a$, $b$ in (3.16) is a highly nontrivial problem).
The work discussed here became Alekseevsky’s dissertation [Ale1892]. It was published as a separate book and is a literal reprint of [Ale1888]-[Ale1889] with of 4 pages added to the beginning. The introduction contains the following sentence:
The known dependence between the Euler function of the second kind [the gamma function] and the sine [i.e., the reflection formula $\Gamma(z)\Gamma(1-z)=\pi/\sin(\pi z)$] made me think to extract from the function $\vartheta_{1}$ of Jacobi a function similar to $1/\Gamma(x)$ and investigate its properties.
By way of analogy with the function $\Gamma(x)$, I came to consider functions whose roots have the form $m\omega+n\omega^{\prime}$, where $m$ and $n$ are negative integers, and $\omega$, $\omega^{\prime}$ are periods of elliptic functions.
As we have seen above, Alekseevsky introduced additional parameters $\omega_{1}$, $\omega_{2}$ to the function $G=G_{2}(x)$. After Barnes [Bar1904], we know that it is possible to introduce similar parameters $\omega_{1}$, …, $\omega_{n}$ to the functions $G_{n}(x)$ and get ‘multigamma functions’. Such functions (we use a non-standard notation) satisfy the collection of identities
$G_{n}(x+\omega_{j};\omega_{1},\dots,\omega_{n})=A_{j}\cdot G_{n-1}(x;\omega_{1},\dots,\widehat{\omega}_{j},\dots,\omega_{n})\,G_{n}(x;\omega_{1},\dots,\omega_{n}),$
where $A_{j}=A_{j}(x,\omega_{1},\dots,\omega_{n})$ are certain elementary factors.
Actually, Appell [App1881] considers more general functions $O(x;\omega,\omega_{1},\dots,\omega_{n})$, which are products of certain exponentials
$\exp\bigl\{a_{0}+a_{1}x+\dots+a_{n}x^{n}\bigr\}$
and expressions
$x\cdot\prod_{w}\Bigl(1-\frac{x}{w}\Bigr)\exp\Bigl\{\sum_{j=1}^{n+1}\frac{1}{j}\Bigl(\frac{x}{w}\Bigr)^{j}\Bigr\},$
where
$w=m\omega+\sum_{j}m_{j}\omega_{j},$
and the integers $m_{j}$ run in the domain
$m\in\mathbb{Z},\qquad m_{j}\geqslant 0,\qquad(m,m_{1},\dots,m_{n})\neq(0,0,\dots,0).$
It is strange that Appell himself did not introduce the double gamma. In any case, Alekseevsky’s response to Appell’s paper was completely natural.
Certainly, Alekseevsky kept in mind the multiple gamma function with several parameters. The introductory part of dissertation [Ale1892] contains the following ’thesis’:
2. Roots of functions similar to $\Gamma$ have the form
$m_{0}+m_{1}a_{1}+\dots+m_{n}a_{n},$
where $m_{k}$ are negative integers, $a_{k}$ are parameters of the function $H$, positive integers or complex quantities.
This ’thesis’ had no continuation in the text of the dissertation.
Visiting Leipzig in 1894, Alekseevsky published a short paper [Ale1894] in German in Leipz. Ber. (communicated by Sophus Lie), and his work apparently became known due to this publication. The paper is a sketch of [Ale1888]-[Ale1889] with some additional claims at the end of the paper. In particular,
Daher folgt:
(1) [^18] These numbers are my addition.Functionen sind construirt aus einer Function $H$, die analog der Function $\Gamma$ ist.
(2) Dieser Satz giebt auch eine ähnliche Beziehung der Modulfunctionen von Hermite zur Function $H$.
(3) Aus der Betrachtung der Functionen höherer Ordnung, die analog der Function $\Gamma(x)$ sind, folgt:
Die Euler’schen Functionen von Herrn Appell sind Producte von Functionen, die analog der Gammafunction sind.
Clearly, in (3) Alekseevsky claims that an Appell function is a product of two multiple gammas [^19] In (1) ’Heine function’ is (3.20); it is the $q$-gamma function up a normalization, which can be expressed in $G(x;a)$ by the reflection formula (see my next subsection). Since we can express theta as a product of 4 double gammas, we can express any double periodic function, see for example [Akh1970], §19.
But I do not understand the meaning of ’Modulfunctionen von Hermite’ and the claim (2).
In any case, the multiple gamma function with several parameters had not appear in Alekseevsky’s published papers and actually was introduced by Barnes [Bar1904]. In Subsect. 5, we continue this discussion.
3.6. The reflection formula. In particular, the Leipzig paper [Ale1894] contains a counterpart of the reflection formula for the gamma function,
$G(1+x;\alpha)\,G(-x;-\alpha)=C\cdot O(x),$
where
(3.20) $O(x):=\prod_{k=1}^{\infty}(1-q^{2k}e^{2\pi ix}),\qquad\text{where $q=\pi i\alpha$,}$
(this follows from the previous subsection’s reasonings). Applying the Jacobi triple identity (see, for example [Akh1970], §18), we can easily express the standard theta functions in the double gamma.
3.7. A counterpart of the $\psi$-function. Let us return to the work [Ale1888]-[Ale1889]. Recall that the classical $\psi$-function (see, for example [Erd1953], 1.7) is defined by
$\psi(z)=\frac{d}{dz}\ln\Gamma(z)=\frac{\Gamma^{\prime}(z)}{\Gamma(z)}=\int_{0}^{\infty}\Bigl(\frac{e^{-t}}{t}-\frac{e^{-zt}}{1-e^{-t}}\Bigr)\,dz.$
In §3 Alekseevsky introduces a similar function
$\varphi(x)=\frac{d}{dx}\ln G(x)$
(in fact, both functions appeared above in (3.17), (3.18)). The function $\varphi$ can be expressed explicitly in terms of $\psi$,
$\varphi(x)=(x-1)(\psi(x)-1)+\varphi(1),\qquad\varphi(1)=-\tfrac{1}{2}+\tfrac{1}{2}\ln 2\pi.$
for $a=1$ this gives (3.6). So, here he has not come to a new special function, but he has found some interesting formulas.
Expanding $e^{ux}$ in (3.4) in $(ux)^{n}$, we can get (§7) an expansion of $\ln G(x+a)$ in a power series,
$\ln G(x+a)=\ln G(a)+x\varphi(a)+\frac{x^{2}}{2}\varphi^{\prime}(a)++\sum_{j\geqslant 3}(-1)^{j+1}\frac{x^{j}}{j}C_{j},$
where
$C_{j}=\sum_{k=1}^{\infty}\frac{k}{(a+k-1)^{j}}.$
For $a=1$ we have $C_{j}=\zeta(j-1)$, for general $a$ we have an expression in terms of Hurwitz zeta $\zeta(a;j)$ and $\zeta(a;j-1)$.
Alekseevsky obtains some other interesting identities for the function $\varphi$, for instance,
$\varphi(1+x)=-\tfrac{1}{2}+\tfrac{1}{2}\ln 2\pi-x(1+\gamma)+\sum_{k=1}^{\infty}\frac{x^{2}}{k(x+k)},$
and
$\varphi(a+x)=\varphi(a)+x\psi(a)+\sum_{j=1}^{\infty}(-1)^{j-1}\frac{x(x-1)\dots(x-j+1)}{j(j+1)(a)_{j}}.$
4. On the history of the double gamma function
4.1. Kinkelin’s work, 1860. Hermann Kinkelin [Kin1860] introduced a function [^20] His original notation is ${\bf G}(x)$. $K(x)$ satisfying the identity
(4.1) $K(x+1)=x^{x}K(x),\qquad K(1)=1$
by (§6)
$\ln K(x)=\int_{0}^{x}\ln\Gamma(x)\,dx+\frac{x(x-1)}{2}+\frac{1}{2}x\ln 2\pi.$
Comparing this expression with (3.9), we see that the double gamma function $G(x)$ is equivalent to $K(x)$, namely
$G(x)\,K(x)=\Gamma(x)^{x-1}$
as it was observed by Beaupain [Bea1903].
Kinkelin obtained some properties of the function $K(x)$. In particular, he wrote its representations as limits of products and expressions of elementary integrals in terms of $K(x)$ (§§10-12). The most remarkable of his identities for $K(x)$ is the multiplication formula, which was established in a better form than Alekseevsky’s (3.8):
(4.2) $K(nx)=n^{\frac{1}{2}nx(nx-1)+\frac{1}{12}}\,\widetilde{\omega}^{-\frac{1}{2}(n^{2}-1)}\cdot\prod_{j=0}^{n-1}K\bigl(x+\tfrac{j}{n}\bigr)^{n},$
where $\widetilde{\omega}$ is a certain constant.
He found several expressions for this constant. For each $n>1$,
(4.3) $\ln\widetilde{\omega}=\frac{n}{n^{2}-1}\Bigl(\frac{1}{12}\frac{\ln n}{n}+\sum_{j=0}^{n-1}\ln K\Bigl(\frac{j}{n}\Bigr)\Bigr),$
see §8. In particular, for $n=2$ this means
$\ln K\bigl(\tfrac{1}{2}\bigr)=\tfrac{3}{4}\ln\widetilde{\omega}-\tfrac{1}{24}\ln 2.$
So, the constant $\widetilde{\omega}$ is connected to $G\bigl(\frac{1}{2}\bigr)$, which appears in many formulas for the double gamma, by a simple relation. Passing to the limit as $n\to\infty$ in (4.3), we get
$\ln\widetilde{\omega}:=2\int_{0}^{1}\ln K(x)\,dx.$
He also expressed $\widetilde{\omega}$ in terms of the zeta-function $\zeta(x)=\sum_{n=1}^{\infty}n^{-x}$,
$\tfrac{1}{2}\ln\widetilde{\omega}=-\tfrac{1}{24}+\frac{1}{3}\gamma+\sum_{\lambda=1}^{\infty}\frac{\zeta(2\lambda+1)-1}{(2\lambda+1)(2\lambda+3)},$
where $\gamma$ is the Euler constant (see §8). Another expression is given by the following counterpart (§9) of the Raabe formula (see [Erd1953], 1.9(19)) for the gamma function:
$\int_{x}^{x+1}K(t)\,dt=\tfrac{1}{2}\ln\widetilde{\omega}+\tfrac{1}{4}x^{2}(2\ln x-1)$
(Alekseevsky later repeated this formula).
Kinkelin also establishes the asymptotics (§8):
$K(n+1)=\prod_{j=1}^{n}j^{j}\sim\widetilde{w}^{1/2}\,e^{-\frac{1}{4}n^{2}+\frac{1}{12}}n^{\frac{1}{2}(n^{2}+n)+\frac{1}{12}},\qquad n\to+\infty.$
The expression $A:=(\widetilde{\omega})^{1/2}e^{1/12}$ is now called the Glaisher–Kinkelin constant, cf. Glaisher [Gla1877].
Kinkelin’s work didn’t attract immediate attention and became a subject of interest only in the second half of the 1890s. The problem (4.1) seems even more artificial than Alekseevsky’s (3.1), so it is natural to ask about his way to the ’Kinkelin function’. Apparently, it was a result of the following ’speleology’.
For a function $f$, denote
$S_{n}(x):=\sum_{j=0}^{n-1}f\Bigl(x+\frac{j}{n}\Bigr).$
In his previous work, Kinkelin [Kin1854] tries to find all functions $f(x)$ satisfying the ’addition theorem’ in the following sense: for each $n>0$ there are constants $\varphi(n)$, $\psi(n)$ such that
(4.4) $S_{n}(x)=\psi(n)\cdot nf(nx)+\varphi(n).$
Apparently, this work was a response to the paper by Joseph Raabe [Raa1851] on Bernoulli polynomials $B_{p}(x)$. Recall that they are defined by the generating function:
(4.5) $\sum_{p=0}^{\infty}B_{p}(x)\frac{t^{p}}{p!}=\frac{te^{xt}}{e^{t}-1}.$
Their main property is the difference equation
$B_{p}(x+1)-B_{p}(x)=px^{p-1}.$
Raabe showed that they satisfy the addition theorem of the type (4.4):
$\sum_{j=0}^{n-1}B_{p}\Bigl(x+\frac{j}{n}\Bigr)=n^{1-p}B_{p}(nx)$
(this easily follows from (4.5)). In fact, Raabe used another normalization of the Bernoulli polynomials, $\widetilde{B}_{p}(x)=B_{p}(x)-B_{p}(0)$, and his $\varphi(n)$ differed from 0).
Kinkelin examines other solutions of the problem (4.4) (I am not sure that his proof is precise). There are also solutions
$f_{q}(x)=\frac{(-1)^{q}}{(q+1)!}\,\frac{d^{q+1}}{dx^{q+1}}\ln\Gamma(x).$
Notice that now $f_{q}(x)$ satisfies the difference equations $f_{q}(x+1)-f_{q}(x)=\frac{1}{x^{q+1}}$. This family also was not ‘new’ in Kinkelin’s time.
However, he tries to vary the condition (4.4). The function $f(x)=\ln\Gamma(x)$ satisfies the addition theorem
(4.6) $\sum_{j=0}^{n}f\bigl(x+\tfrac{j}{n}\bigr)=f(nx)+\frac{n-1}{2}\ln(2\pi)+\bigl(\tfrac{1}{2}-nx\bigr)\ln n,$
which is equivalent to the multiplication formula for the gamma function. Also, it satisfies the difference equation
(4.7) $f(x+1)-f(x)=\ln x.$
We take primitives from both sides of identities (4.6), (4.7) and get two equations for the primitive $F(x):=\int f(x)\,dx$. Equation (4.7) yields
$F(x+1)-F(x)=x\ln x-x+C.$
Correcting $F(x)\mapsto\widetilde{F}(x)$ by an appropriate quadratic summand, we can reduce the last equation to the form
$\widetilde{F}(x+1)-\widetilde{F}(x)=x\ln x,$
i.e., the equation (4.1) for $\ln K(x)$. The integration of (4.6) gives an addition theorem for $\widetilde{F}(x)$ or a multiplication theorem for the Kinkelin function $K(x)$.
4.2. The double gamma function, 1899-1907. The work of Alekseevsky became known (perhaps due to the publication [Ale1894]). After this, the double gamma function was a topic of publications by Ernest Barnes [Bar1899], [Bar1900.1], [Bar1900.2], [Bar1901], [Bar1903], [Bar1904], Jean Beaupain [Bea1903], [Bea1904], [Bea1907], Godfrey Hardy [Har1905.1], [Har1905.2], and Vladimir Steklov [Ste1904]. The most important result of these works was Barnes’s introduction of the multiple gamma functions [Bar1904].
In particular, the third Barnes’s publication [Bar1899] and his further works [Bar1900.1][Bar1903] were devoted to the double gamma. In [Bar1903] he found the first application: he expressed the electrostatic capacity coefficients of a pair of non-intersecting spheres in terms of the double gamma.
In the work [Bar1904] he introduced a new topic, namely the multiple gamma function; it is a function $G_{n}(x)$ depending on parameters $\omega_{1}$, …, $\omega_{n}$ and satisfying functional equations
$G_{n}(x+\omega_{j};\omega_{1},\dots,\omega_{n})=A_{j}\cdot G_{n-1}(x;\omega_{1},\dots,\widehat{\omega}_{j},\dots,\omega_{n})\,G_{n}(x;\omega_{1},\dots,\omega_{n}),$
where $A_{n}$ are certain elementary expressions and $G_{n-1}$ is the multiple gamma of the previous level.
Barnes became famous due to his later works of 1907–08 on the ’Mellin–Barnes integrals’, i.e., integrals of the type
$\int_{-i\infty}^{\infty}z^{-s}\frac{\prod\Gamma(a_{k}+s)\prod\Gamma(b_{l}-s)}{\prod\Gamma(c_{m}+s)\prod\Gamma(d_{n}-s)}\,ds.$
Integrals of such kind appeared earlier in works by Salvatore Pincherle and Hjalmar Mellin. In any case, Barnes’s research was a breakthrough — thanks to him such integrals became a powerful and flexible tool of theory of special functions, see for example [Whi1954], [AAR1999], [Mari1978] (in particular, Mellin–Barnes integrals were the main tool of creations of tables of integrals [PBM1985], see [PBM1989]; see also the impressive table of Mellin-Barnes integrals in volume 3 of [PBM1985], Chapter 8).
4.3. The higher Kinkelin functions. Beaupain [Bea1903] introduced higher Kinkelin functions, i.e., functions, which satisfy the identity
$K_{n}(x+1)=x^{x^{n}}K_{n}(x),\qquad K_{n}(1)=1.$
Alekseevsky [Ale1904] showed that systems of functions $G_{n}(x)$ (see above Subsect. 3) and $K_{n}(x)$ can be expressed through each other,
(4.8) $\ln K_{n}(x)=\sum_{j=0}^{n}(-1)^{j}(\Delta^{j}x^{n})\,\ln G_{j}(x+j),$
where $\Delta f(x):=f(x+1)-f(x)$.
4.4. The gap, 1908-1975. After 1907, the double gamma was forgotten.
As was observed in [Ada2014], this function was mentioned in the second edition [WhWa1912] of E. T. Whittaker, G. N. Watson A course of Modern Analysis (Chapter XII, Examples 48-50, namely (3.6) as a definition, (3.1), (3.5), (3.11)). Next, it was mentioned in the fourth edition [Grad1963] of Gradshteyn, Ryzhik Tables of integrals, sums, and series, lines 6.441, 8.333, with references to Whittaker–Watson.
The double gamma was absent in other handbooks on special functions as Higher transcendental functions [Erd1953], or Prudnikov, Brychkov, Marichev, [PBM1985].
I have seen (mainly, due to [Gus1952]) a few other appearances of such functions in literature. In 1933, L. Bendersky [Ben1933] published a large detailed investigation on higher Kinkelin functions (he probably didn’t know about his predecessors).
Boris Rimsky-Korsakov published a paper [Rim1947] on the reciprocity theorem for higher Kinkelin functions:
$\frac{\Bigl\{\prod\limits_{k=0}^{q-1}K_{n}\bigl(pz+\frac{kp}{q}\bigr)\Bigr\}^{q^{n}}}{\Bigl\{\prod\limits_{k=0}^{p-1}K_{n}\bigl(qz+\frac{kq}{p}\bigr)\Bigr\}^{p^{n}}}=\Bigl(\frac{p}{q}\Bigr)^{\frac{p^{n}}{n+1}\sum_{k=0}^{n-1}B_{n+1}\bigl(qz+\frac{kq}{p}\bigr)}\pi_{n}^{(q^{n+1}-p^{n+1})/2},$
where $B_{m}(\cdot)$ are the Bernulli polynomials and $\pi_{n}$ are certain constants. For $n=1$ this statement is the Schobloch reciprocity formula ([Scho1884], [Srin2007]) for the gamma function. He also published the paper [Rim1957] and PhD thesis [^21] I have not seen it. [Rim1957.1] about further generalizations of the gamma function.
The double gamma function was discussed in the obituary of Barnes, written by Edmund Whittaker [Whi1954].
4.5. A new appearance. The double gamma function appeared again in the work of Takuro Shintani [Shi1976]-[Shi1977]. After that, the function started showing up in various math problems, I only present a few references [Vig1979], [Sar1987], [Var1988], [NiYo2009], [Ada2014], [Ost2013], [AlKu2023], [KarKu2024].
Shintani [Shi1977] also introduced the double sine
$\mathcal{S}_{2}(x;\omega_{1},\omega_{2})=(2\pi)^{\frac{\omega_{1}+\omega_{2}-2x}{2\omega_{1}}}\omega_{2}^{\frac{x(\omega_{1}+\omega_{2})}{\omega_{1}\omega_{2}}}\frac{G\bigl(\frac{x}{\omega_{1}};\frac{\omega_{2}}{\omega_{1}}\bigr)}{G\bigl(1-x+\frac{\omega_{2}}{\omega_{1}};\frac{\omega_{2}}{\omega_{1}}\bigr)}=\\ =\exp\Bigl\{\int_{0}^{\infty}\Bigl(\frac{\sinh\bigl(z-\frac{1}{2}(\omega_{1}+\omega_{2})\bigr)}{\sinh\bigl(\frac{1}{2}\omega_{1}t\bigr)\,\sinh\bigl(\frac{1}{2}\omega_{1}t\bigr)}-\frac{1}{\omega_{1}\omega_{2}t}(2z-\omega_{1}-\omega_{2})\Bigr)\frac{dt}{t}\Bigr\}$
(a special case of the double sine corresponding to equal periods was considered by Otto Hölder [Hol1886]). This function is important by itself. Much later (and independently) Ludvig Faddeev [Fad1995] introduced the quantum dilogarithm, which is very similar to the double sine. However, it appeared in a completely new context, and the term ’quantum dilogarithm’ is now even more common than ’double sine’.
For instance, see some nice applications of the double sine in representation-theoretical spirit (which were the reason of my interest in this topic) in [Kash1998], [FoCh1999], [KLS2002]).
4.6. Why were the double gamma and the multiple gamma forgotten 70 years? I’ll try to speculate on possible causes. One of the reasons is clear: the function was different significantly from special functions that were considered at that time. In my opinion, there were also some accidental reasons.
Firstly, both Alekseevsky and Barnes (1874-1953) left mathematics. Alekseevsky became a professor of mechanics and a rector of a technological institute (his last research publication was in 1904). Barnes became an Anglicanic deacon, a theologist, and later a bishop of Birmingham, two of his last research publications are dated 1908 and 1910, see [Whi1954]).
Secondly, this can be partially related to the structure of [Ale1888]-[Ale1889]. The work can seem like a free fantasy, since its real external relations become clear only at the end of the second paper. There was a chance for a new appearance of the double gamma after a historical-mathematical paper by Viktor Gussov [Gus1952], which contained a good exposition of [Ale1888]-[Ale1889] (I think many Russian mathematicians have seen this article and were surprised reading it). But that didn’t happen.
Thirdly, a meticulous reader could decide that the actual topic of the works on the double gamma are the functions $\ln G(x)$, $\ln G_{n}(x)$, $\ln G(x;a)$. For instance, in Alekseevsky’s papers, the double gamma itself (without a logarithm) never appears in the right hand sides of integrals, which contrasts with the behavior of $\Gamma(x)$. The classical gamma function appears in the right hand sides of a huge number of definite integrals, see for example [Grad1963], [PBM1985]; on the other hand, appearances of $\ln\Gamma(\cdot)$ in the right hand sides of integrals are relatively rare. Next, integrals like (3.10), (3.11) are not too impressive (the right hand side is not better than the left hand side) and admit other transformations. Only in §22 of [Ale1888][Ale1889] a reader can understand that products of the double gammas can participate in Heine’s theory of $q$-hypergeometric functions [^22] Also, $q$-hypergeometric functions were almost unknown in Russia until the appearance of interest in representations of classical groups over finite and $p$-adic fields in the 1960-s. So, to a Russian reader, motivation of [Ale1888]-[Ale1889] was unclear.,
Now there are no reasons for such doubts. Ludvig Faddeev [Fad1995] observed that the double sine is a kernel of a natural integral operator in $L^{2}(\mathbb{R})$ satisfying highly nontrivial identities, see [Fock1997], [Kash1998], [FoCh1999]. See also the work [KLS2002], where certain spectral measures for quantum groups are expressed as products of the double sines. Recall that products of the usual gamma functions appear as spectral densities for differential and difference operators (as far as they admit explicit evaluations), such products also appear as spectral densities in numerous problems of harmonic analysis for semisimple Lie groups.
See, also a recent work [KarKu2024].
The term ’Alexeiewsky function’ was used in several papers published in 1899-1907. The widely used modern term ’Barnes double gamma’ was originated by Shintani [Shi1976][Shi1977], who found old Barnes works. Alekseevsky, who was mainly the author of one big work [Ale1888], [Ale1889], [Ale1894], [Ale1904], was forgotten at that time, while Barnes was well known due to the Barnes integrals (see above).
5. Alekseevsky: his historiography, lost work, and publications
5.1. Historial sources and historiography. I used the following contemporary historical sources on the biography of Alekseevsky:
— Obituary written by the president (1906-1946) of the Kharkov Mathematical Society Dmitry Sintsov [Sin1917].
— Steklov and Sintsov’s review [StSi1904] of Alekseevsky’s works (apparently written by Steklov) supporting Alekseevsky’s application to the position of extra-ordinary professor in Kharkov University.
— Steklov’s correspondence published in [Ste1991].
As told above, Alekseevsky’s work was forgotten. But his name several times was mentioned in histories of mathematics and education in Russia.
— Viktor Gussov wrote an interesting and thorough dissertation [Gus1950] on investigations of the gamma function and the Bessel functions in Russia (the work was published in [Gus1952][Gus1953]). He found Alekseevsky’s works and exposed them in [Gus1950], [Gus1952]. Aleksey Markushevich was an opponent of Gussov’s thesis and used it in his ’Essays’ on the history of complex analysis [Mark1951]. However, he did not mention Alekseevsky in his ’Essays’, nor in his historical investigation [Mark1981], nor in his famous treatise Theory of functions of complex variable. Clearly, Gussov’s paper [Gus1952] was used in the publications mentioned below.
— Some information is contained in Èsfir Bakhmutskaya’s work [Bakh1967] on mathematics in Kharkov University.
— The Tomsk period of Alekseevsky’s biography was well-investigated in several works about the history of science in Tomsk (see Nikolay Krulikovsky [Kru1967], [Kru1967.1], Yulia Granina [Gran2003], Belomestnyh [Bel2000]).
5.2. The multiple gamma function and a lost work. Let us return to the discussion in Subsect. 3. In their review of Alekseevsky works [StSi1904], Steklov and Sintsov wrote:
Now as well, Alekseevsky continues to successfully develop his theory of functions, which he now calls gammamorphic. He came to functions of an even more general type, for which Kinkelin functions [Steklov keeps in mind $G_{n}(x)$] are limit cases, and which are related to the known Riemann $\zeta(s)$ and many other functions important in Analysis and its applications.
These recent studies were reported at the meetings of the Kharkov Mathematical Society in 1902 and 1903. Part of them are currently beeing printing in our society[s Communications.
We see in [Khar1904] the following titles of Alekseevsky’s talks at the Kharkov Mathematical Society:
15.02.1902. On the Riemann $\zeta$-function of three arguments.
03.12.1902. On the Riemann $\zeta$-function of two arguments.
24.10.1903. Note on the Kinkelin functions.
31.10.1903. Towards the theory of gammamorphic functions.
Clearly, the third talk corresponds to [Ale1904].
Sintsov [Sin1917] mentions that an article ‘On Riemann function $\zeta$’ disappeared and writes that there is hope to find it in Alekseevsky’s posthumous papers.
It seems clear that the topic of these works were the multiple gamma functions and related $\zeta$-functions of the type
$\sum_{m,n\geqslant 0}(z+n\omega_{1}+m\omega_{2})^{-s}$
(something similar to Barnes [Bar1904]).
In any case, Alekseevsky has not published anything on this topic except the claims in [Ale1889], [Ale1892], [Ale1894] quoted above [^23] Letter of Steklov to Lyapunov, 29.12.1902: Alekseevsky, as usual, drags on and does not submit his article [Yu. N.: clearly, this sentence refers to [Ale1904]], although he says that he will submit it ‘one of these days’, but this ‘one of these days’ has been going on since September..
5.3. Alekseevsky’s publications. The list of his research publications is contained in our bibliography, [Ale1884.1][Ale1904] (all of them belong to the Kharkov period of his work).
Russian library catalogues show several editions of his lecture courses and problem books (in Kharkov Technological Institute and in Tomsk Technological Institute):
Алексеевский В. П. Дифференциальное исчисление : Лекции, чит. в весен. семестре 1900 г. Харьков: типо-лит. С. Иванченко, 1900., 317 с.
Задачи по дифференциальному и интегральному исчислениям, предлагавшиеся проф. Алексеевским в Х[арьковск.] Т[ехнол.] И[н-те]. - Харьков : изд. студ.-технологов Н. Александровича и Н. Праведникова, 1907. - [2], 59 с.
Динамика сост. по лекциям и под ред. В. П. Алексеевского студ. И. Никишовым, 1912-1913 уч. г. ; Том. технологич. ин-т императора Николая II. Типо-лит. Сиб. Т-ва Печатн. Дела 1913.
Сборник задач по динамике с подробными решениями, предлагавшихся профессором В. П. Алексеевским в Томском технологическом институте в 1913-14 уч. году на III и IV сем. Сост. Г. Берштейн, студент Мех. отд-ния. - Томск: типо-лит. Сиб. т-ва печ. дела, 1914 , 47 p.
Алексеевский В.П. Динамика : Лекции 1915-1916 уч. г. Том. технол. ин-т. Томск: Т-во "Печатня С. П. Яковлева 1916, 237 с.
Also, [Sin1917] and [Gran2003] contains information about some additional lithographic editions of his lecture courses in Kharkov University and in Tomsk Technological Institute.
6. Conclusion
The double gamma was introduced in the work of Vladimir Alekseevsky in 1888-1889, he also derived main identities involving this function (cf. the list of identities [AlKu2023]). His publication [Ale1894] had a big influence on the early works by Ernest Barnes, who also introduced and investigated the multiple gamma function [Bar1904]. The double gamma function was Alekseevsky’s main work and he was forgotten for the same period that the double gamma itself was forgotten.
In this paper, we observed an interesting case: that the results of a work, which were very difficult to replicate, were forgotten for 70 years. This happened despite the fact that some of the authors writing about the double gamma, were well-known and that the results were vivid. The cause was probably an unfortunate confluence of several circumstances.
Modern mathematicians use complicated languages, having trouble understanding each other. Large losses of nontrivial information over the intervals of several decades are unavoidable (and is, in fact, already occuring). In our case, however, the papers on special functions and complex analysis of 1880-1910s were written in a simple language one accessible for mathematicians of 1970-s and since. They could have been easily read and used.
Список литературы
  • [Ada2014] Adamchik V.S. Contributions to the Theory of the Barnes Function, Int. J. Math. Comput. Sci. 9 (2014), no. 1, 11-30.
  • [Akh1970] Akhiezer N. I. Elements of the theory of elliptic functions. American Mathematical Society, Providence, RI, 1990.
  • [Ale1884.1] Алексеевский В. П. Об интегрировании уравнения $\frac{d^{n}y}{dz^{n}}+\frac{\alpha}{z}\frac{d^{n-1}y}{dz^{n-1}}+\beta{y}=0,$ (Russian) Сообщ. и протоколы заседаний Матем. общ. при Император. Харьков. унив. 1884 года, 1884, 1, 41-64; [Alexeiewsky V. P. On the integration of the equation $\frac{d^{n}y}{dz^{n}}+\frac{\alpha}{z}\frac{d^{n-1}y}{dz^{n-1}}+\beta{y}=0,$ Charkow Ges., 1884, 1, 41-64] JFM [^24] JFM database; the electronic version of Jahrbuch über die Fortschritte der Mathematik, 1868-1942. It is contained in the Zentralblatt für Mathematik database. 16.0290.01.
  • [Ale1884.2] Алексeевский В. П. Заметка об обобщении уравнения Риккати. (Russian) Сообщ. и протоколы заседаний Матем. общ. при Император. Харьков. унив. 1884 года, 1884, 1, 80-82 [Alexeiewsky V. P. Note on the generalization of the Riccati equation. [Charkow Ges.], 1884, 1, 80-82] JFM 16.0290.01]
  • [Ale1884.3] Алексеевский В. П. Об интегрировании уравнения $x^{2}y^{\prime\prime\prime}+Axy^{\prime\prime\prime}+By^{\prime}+Cx^{\mu}y=0,$ (Russian) Сообщ. и протоколы заседаний Матем. общ. при Император. Харьков. унив. 1883 года, 1884, 2, 115-126 (Russian) [Alexeiewsky W. On the integration of the equation $x^{2}y^{\prime\prime\prime}+Axy^{\prime\prime\prime}+By^{\prime}+Cx^{\mu}y=0,$ Charkow Ges., 1884, 2, 115-126] JFM 16.0288.01.
  • [Ale1885] Алексеевский В. П., Об интегрировании одного линейного диффенциального уравнения nn-го порядка (Russian) Сообщ. и протоколы заседаний Матем. общ. при Император. Харьков. унив. 1884 года, 1885, 3, 222-232. [Alexeiewsky V. P. On integration of one linear differential equation of $n$-th order, Charkow Ges., 1885, 3, 222-232] JFM 17.0328.01
  • [Ale1888] Алексеевский В. П. О функциях подобных функции гамма (Russian), Сообщ. Харьков. матем. общ. Вторая сер., 1:1 (1888), 169-192 [Alexeiewsky W. P. (Alexejewsky W.) On the functions that are similar to the gamma function.], Charkow Ges., 1:1 (1888), 169-192] JFM 22.0439.03.
  • [Ale1889] Алексеевский В. П. О функциях подобных функции гамма (Russian), Сообщ. Харьков. матем. общ. Вторая сер., 1:2 (1889), 193-238. [Alexeiewsky W. P. (Alexejewsky W.) On the functions that are similar to the gamma function., Charkow Ges., 1:2 (1889), 193-238] JFM 25.0751.02
  • [Ale1892] Алексеевский В. П. О функциях подобных функции Гамма. (Russian) Диссертация. Харьков: тип. Зильберберга, 1892, 70 с. [Alexeevsky V.P. On the functions that are similar to the gamma function. Dissertation. Printing house of Zilberberg, Kharkov, 1892, 70p.]
  • [Ale1894] Alexejewsky W. P. Ueber eine Classe von Funktionen, die der Gammafunktion analog sind (German), Berichte über die Verhandlungen der Königlich-Sächsischen Gesellschaft der Wissenschaften zu Leipzig, 46 (1894) 268-275. JFM 25.0751.02
  • [Ale1895] Алексеевский В. П. Об автоморфной функции аналогичной экспонентной (Russian), Сообщ. Харьков. матем. общ. Вторая сер., 4 (1895), 253-262 [Alexeiewsky V. P. On an automorphic function that is analogous to the exponential, Charkow Ges., 4 (1895), 253-262] JFM 26.0466.02
  • [Ale1899.1] Алексеевский В. П. О законе взаимности простых чисел (Russian), Сообщ. Харьков. матем. общ. Вторая сер., 6 (1899), 200-202 [Alexejevsky W. P. On the reciprocity law of primes, Charkow Ges., 6 (1899), 200-202] JFM 30.0184.03
  • [Ale1899.2] Алексеевский В. П. Об определении длины в неэвклидовой геометрии (Russian) Сообщ. Харьков. матем. общ. Вторая сер., 6 (1899), 139-153 [Alexejewsky W. P. On a definition of a length in the non-Euclidean geometry, Charkow Ges., 6 (1899), 139-153] JFM 29.0411.03
  • [Ale1904] Алексеевский В. П. Зависимость между Кинкелиновыми и гаммаморфными функциями. (Russian) Сообщ. Харьков. матем. общ. Вторая сер., 8 (1904), 123-135. [Alexejewsky W. P. Dependence between Kinkelin and gamma-morphic functions, Charkow Ges., 8 (1904), 123-135. JFM 34.0490.01
  • [AlKu2023] Alexanian S., Kuznetsov A. On the Barnes double gamma function. Integral Transforms Spec. Funct. 34 (2023), no. 12, 891-914.
  • [AAR1999] Andrews G. E., Askey R., Roy R. Special functions. Cambridge: Cambridge University Press., (1999).
  • [App1881] Appell P. Sur une classe de fonctions analogues aux fonctions Eulériennes. (French) Klein Ann. XIX, 84-102 (1881), JFM 13.0388.01.
  • [Bakh1967] Bakhmutskaya È. Ya. Mathematics in Kharkov university. Kharkov mathematical society. (Russian) In History of national mathematics in four volumes. Vol. 2, 1801-1917. (eds. Z. Shtokalo, A. N. Bogoljubov, Ju. A. Mitropol’skiĭ, I. B. Pogrebysskiĭ, E. Ja. Remez, K. A. Rybnikov, Ju. D. Sokolov and V. S. Sologub), 460-472, Naukova Dumka, Kiev, 1967.
  • [Bar1899] Barnes E. W. The Genesis of the Double Gamma Functions Proc. Lond. Math. Soc. 31 (1899), 358-381. JFM 30.0389.03
  • [Bar1900.1] Barnes E. W. The theory of the $G$-function. Quart. J. 31, 264-314 (1900). JFM 30.0389.02
  • [Bar1900.2] Barnes E. W. The theory of the double gamma Function. JFM 31.0441.01 Lond. R. S. Proc. 66, 265-266 (1900).
  • [Bar1901] Barnes E. W. The theory of the double gamma function. Lond. Phil. Trans. (A) 196, 265-387 (1901). JFM 32.0442.02
  • [Bar1903] Barnes E. W. On the coefficients of capacity of two spheres. Quart. J. 35, 155-175 (1903). JFM 34.0913.01
  • [Bar1904] Barnes E. W. , On the theory of the multiple gamma function, Trans. Camb. Philos. Soc., 19 (1904), 374-425. JFM 35.0462.01
  • [Bea1903] Beaupain J. Sur les fonctions d’ordre supérieur de Kinkelin. (French) JFM 34.0492.06 Belg. Mém. C. et. sav. étr. 59, 67 S. (1903). JFM 34.0492.06
  • [Bea1904] Beaupain J. Sur la fonction $\log\Gamma(a)$. Sur la fonction $\log G_{\lambda}(a)$. (French) 13 S. 16 S. Belg. Bull. Science 62 (1904). JFM 35.0461.03
  • [Bea1907] Beaupain J. Sur la fonction gamma double. (French) Belg. Mém. et. sav. étr. in $4^{\circ}$ (2) 1, 25 S. (1907). JFM 38.0469.04
  • [Bel2000] Belomestnyh V. N., Belomestnyh L. A., Physical-mathematical education in high technical school in Siberia. Part 1. (Russian) Tomsk Polytechnical Univ., 2000.
  • [Ben1933] Bendersky L. Sur la fonction gamma généralisée. (French) Acta Math. 61 (1933), 263-322 , JFM 59.0373.02.
  • [Bou1960] Bourbaki N. Éléments d’histoire des mathématiques. (French) Histoire de la Pensée, IV Hermann, Paris, 1960.
  • [Erd1953] Erdélyi A., Magnus W., Oberhettinger F., Tricomi F. G. Higher transcendental functions. Vols. I, II Based, in part, on notes left by Harry Bateman McGraw-Hill Book Co., Inc., New York-Toronto-London, 1953. Vol. III, 1955.
  • [Fad1995] Faddeev L. Discrete Heisenberg-Weyl group and modular group, Lett. Math. Phys. 34, (1995), 249.
  • [Fock1997] Fock V. V. Dual Teichmüller spaces. Preprint
    . https://arxiv.org/pdf/dg-ga/9702018v1.pdf
  • [FoCh1999] Fock V. V., Chekhov L. O. Quantum Teichmüller spaces. Theoret. and Math. Phys. 120 (1999), no.3, 1245-1259.
  • [Gal2004] Galanova R. A. His thoughts and deeds were devoted to Siberia: about the first rector of TTI (TPU) [^25] Tomsk Techological Institute (TTI) changed couple of names, since 1991 it is Tomsk Technological University, (TTU). E. L. Zubashev. (Russian). Tomsky polytehnik, 2004, 10, 160-163.
  • [Gal2009] Galanova R. A. Efim Luk’yanovich Zubashev: a biographical essay. (Russian) Tomsk Polytechnical University, Tomsk, 2009.
  • [Gla1877] Glaisher, J. W. L. On the product $1^{1}\cdot 2^{2}\cdot 3^{3}\dots n^{n}$. Messenger (2) VII. 43-47 (1877). JFM 09.0190.01
  • [Grad1963] Gradshteyn I. S., Ryzhik I. M. Table of integrals, series, and products. (Russian) Fourth edition prepared by Ju. V. Geronimus and M. Ju. Ceĭtlin. Fizmatgiz, 1963. English transl.: Academic Press, New York-London, 1965.
  • [Gran2003] Granina Yu. G. The professor-mathematician, the second director of TTI V. P. Alekseevsky (Russian), Izvestiya Tomskogo polytechnicheskogo universiteta, [Bulletin of Tomsk polytechnical university], 306, no. 3., (2003) 162-167.
  • [Gus1950] Gussov V. V. From the history of transcendental functions in Russia and USSR. (Russian) Thesis (Ph.D.) – Moscow State University, 1950, 385pp.
  • [Gus1952] Gussov V. V. Works of Russian scholars on the theory of the gamma function.(Russian) Istor.-Mat. Issled. 5 (1952), 421-472.
  • [Gus1953] Gussov, V. V. Development of the theory of cylinder functions in Russia and the USSR. (Russian) Istor.-Mat. Issled. (1953), 355-475.
  • [Har1905.1] Hardy G. H. On the expression of the double zeta-function and double gamma-function in terms of elliptic functions. Cambr. Phil. Trans. 20, 1-35 (1905). JFM 36.0508.03
  • [Har1905.2] Hardy G. H. On double Fourier series, and especially those which represent the double zeta-function with real and incommensurable parameters. Quart. J. 37, 53-79 (1905). JFM 36.0501.02.
  • [Hei1847] Heine E. Untersuchungen über die Reihe
    $1+\frac{(1-q^{\alpha})(1-q^{\beta})}{(1-q)(1-q^{\gamma})}\cdot x+\frac{(1-q^{\alpha})(1-q^{\alpha+1})(1-q^{\beta})(1+q^{\beta+1})}{(1-q)(1-q^{2})}{(1-q^{\gamma})(1+q^{\gamma+1})}\cdot x^{2}+\dots$
    (German) J. Reine Angew. Math. 34, 285-328 ( 1847), ERAM [^26] Electronic Research Archive for Mathematics. 034.0971cj.
  • [Hei1878] Heine E. Handbuch der Kugelfunctionen. Theorie und Anwendungen. (German) Berlin, G. Reimer, 1878. Reprinted Physica-Verlag, Würzburg, 1961.
  • [Hol1886] Hölder O. Ueber eine transcendente Function. (German) Gött. Nachr. 1886, 514-522 (1886). JFM 18.0376.01
  • [KarKu2024] Karp D., Kuznetsov A. Extending the Meijer G-function. Preprint
    . https://arxiv.org/abs/2403.05708.
  • [Kash1998] Kashaev R. M. Quantization of Teichmüller spaces and the quantum dilogarithm. Math. Phys. 43 (1998), no.2, 105-115.
  • [KashN2011] Kashaev R. M., Nakanishi T. Classical and quantum dilogarithm identities. SIGMA Symmetry Integrability Geom. Methods Appl. 7 (2011), Paper 102, 29 pp.
  • [KLS2002] Kharchev S., Lebedev D., Semenov-Tian-Shansky M. Unitary representations of $U_{q}(\mathfrak{s}\mathfrak{l}(2,\mathbb{R}))$, the modular double and the multiparticle $q$-deformed Toda chains. Comm. Math. Phys. 225 (2002), no.3, 573-609.
  • [Khar1904] Извлечение из протоколов заседаний (Russian), Сообщ. Харьков. матем. общ. Вторая сер., 8 (1904), 283-287 [Extracts from the minutes of meetings, Charkow Ges., 8 (1904), 283-287 ]
  • [Kin1854] Kinkelin H. Untersuchung über die Formel
    $nF(nx)=f(x)+f\bigl(x+\tfrac{1}{n}\bigr)+f\bigl(x+\tfrac{2}{n}\bigr)+\dots f\bigl(x+\tfrac{n-1}{n}\bigr).$
    (German), Archiv der Mathematik und Physik, 22 (1854), 189-224.
  • [Kin1860] Kinkelin H. Ueber eine mit der Gammafunction verwandte Transcendente und deren Anwendung auf die Integralrechung. (German) J. Reine Angew. Math. 57, 122-138 (1860). ERAM 057.1509cj
  • [Kru1967] Krulikovskiĭ. N. N. History of the development of mathematics in Tomsk. (Russian) Izdat. Tomsk. Univ., Tomsk, 1967; Second edition 2006.
  • [Kru1967.1] Krulikovskiĭ. N. N. Mathematics in Tomsk Technological Institute. (Russian) In History of national mathematics in four volumes. Vol. 2, 1801-1917. (eds. Z. Shtokalo, A. N. Bogoljubov, Ju. A. Mitropol’skiĭ, I. B. Pogrebysskiĭ, E. Ja. Remez, K. A. Rybnikov, Ju. D. Sokolov and V. S. Sologub), 524-527. Naukova Dumka, Kiev, 1967.
  • [Mari1978] Marichev O. I The method of evaluation of integrals of special functions (theory and tables of formulas) (Russian), Nauka i Tehnika, Minsk, 1978; English version: Marichev O. I. Handbook of integral transforms of higher transcendental functions: theory and algorithmic tables. John Wiley, New York, 1983.
  • [Mark1951] Markushevich A. I. Essays on the history of the theory of analytic functions (Russian). Gosudarstv. Izdat. Tehn.-Teor. Lit., Moscow-Leningrad, 1951.
  • [Mark1981] Markushevich A. I. Theory of analytic functions. (Russian). In Mathematics of the XIXth century, Geometry, Theory of analytic functions (eds. A. N. Kolmogorov, A. P. Yushkevich), Moscow 1981, 115-269 (1981); English transl.: Laptev B. L., Rozenfeld B. A., Markushevich A. I. Mathematics of the 19th century. Geometry, analytic function theory. Birkhäuser, Basel, 1996.
  • [Mol1985] Molin F. È. [Molien Th.] Number systems. (Russian) Translated from the German by L. A. Bokut’, N. N. Krulikovskiĭ. and I. V. L’vov. Edited by A. I. Kostrikin. With appendices by Krulikovskiĭ, Bokut’. and L’vov. Nauka, Sibirsk. Otdel., Novosibirsk, 1985. 126 pp. Zbl [^27] The database Zentralblatt für Mathematik. 0607.01025
  • [NiYo2009] Nikeghbali A., Yor M. The Barnes $G$-function and its relations with sums and products of generalized Gamma convolution variables. Electronic Communications in Probability. 2009; 14:396.
  • [Ost2013] Ostrovsky D. Theory of Barnes beta distributions. Electron. Commun. Probab. 18, Paper No. 59, 16 p. (2013).
  • [PBM1985] Prudnikov A. P., Brychkov Yu. A., Marichev O. I. Integrals and series. Vol. 1-3: Gordon & Breach 1986-1990.
  • [PBM1989] Prudnikov A. P., Brychkov Yu. A., Marichev O. I. Calculation of integrals and the Mellin transform. (Russian) Mathematical analysis, Vol. 27, 3-146, Progress in Science and Technology, VINITI, 1989; Translated in J. Soviet Math. 54 (1991), no. 6, 1239-1341.
  • [Raa1851] Raabe J. L. Zurückführung einiger Summen und bestimmten Integrale auf die Jacob-Bernoullische Function. (German) J. Reine Angew. Math., 1851(42), 348-367. ERAM 042.1172cj
  • [Rim1947] Rimskiĭ-Korsakov B. S. Notes on generalized multiplication theorems of Bernulli polynomials and Kinkelin functions. (Russian) Trudy MAI [Moscow Aviation Institute], 1, 1947
  • [Rim1957] Rimskiĭ-Korsakov B. S. A version of the construction of a theory of generalized gamma-functions based on the Laplace transform. Amer. Math. Soc. Transl. (2)17(1961), 201-217; Translated from: Moskov. Oblast. Ped. Inst. Uch. Zap. [Московский Областной Педагогический Институт. Ученые записки.] 57 (1957), 121-141.
  • [Rim1957.1] Rimskiĭ-Korsakov B. S. An investigation on theory of generalizations of the Gamma-function (Russian). Dissertation for the degree of candidate of physical and mathematical sciences, Moscow, 1957.
  • [Sar1987] Sarnak P. Determinants of Laplacians. Comm. Math. Phys.110 (1987), no.1, 113-120.
  • [Scho1884] Schobloch, J. A. Über Beta- und Gammafunctionen. (German) Halle. Nebert (1884). JFM 16.0395.01.
  • [Shi1976] Shintani T. On Kronecker limit formula for real quadratic fields. Proc. Japan Acad.52 (1976), no.7, 355-358.
  • [Shi1977] Shintani T., On a Kronecker limit formula for real quadratic fields, J. Fac. Sci. Univ. Tokyo, Sect. 1A 24 (1977), 167-199
  • [Sin1917] Синцов Д. М. В. П. Алексеевский (1858-1916). (Russian) Сообщ. Харьков. матем. общ. Вторая сер., 15:5-6 (1917), 288-295 [Sintsov D. M. V. P. Alexeevsky (1858-1916). Charkow Ges., 15, no. 5-6 (1917), 288-295].
  • [Srin2007] Srinivasan G. K. The gamma function: an eclectic tour. Amer. Math. Monthly 114 (2007), no. 4, 297-315.
  • [Ste1904] Stekloff W. A. [Steklov V. A.] Remarques relatives aux formules sommatoires d’Euler et de Bool (French), Сообщ. Харьков. матем. общ. Вторая сер. [Charkow Ges.], 8 (1904), 136-195. JFM 35.0263.01
  • [Ste1991] Steklov V. A. Correspondence with national mathematicians. Recollections. (Russian), eds. V. S. Vladimirov, E. P. Ozhigova, V. S. Sobolev, Leningrad, Nauka, 1991.
  • [StSi1904] Стеклов В. А., Синцов Д. М. Отзыв профессоров Стеклова и Синцова об ученых трудах прив.-доцента В. П. Алексеевского. (Russian) Ученые записки Императорского Харьковского университета, 1904, 1, 1-11 [Review of professors Steklov and Sintsov on the scientific works of privatdozent V. P. Alekseevsky. Scientific notes of the Imperator’s Kharkov University, 1904, 1, 1-11]
  • [Var1988] Vardi I. Determinants of Laplacians and Multiple Gamma Functions. SIAM Journal on Mathematical Analysis, 19(2) (1988), 493-507.
  • [Vig1979] Vignéras M. F. L’équation fonctionelle de la fonction zêta de Selberg du groupe modulaire $SL(2,\mathbb{Z})$ (French), Astérisque 61 (1979), 235-249
  • [Wae1985] van der Waerden B. L. A history of algebra. From al-Khwārizmī. to Emmy Noether. Springer-Verlag, Berlin, 1985.
  • [Whi1954] Whittaker E. T. Ernest William Barnes, 1874-1953. Obit. Notices Roy. Soc. London 9 (1954), 15-25.
  • [WhWa1912] Whittaker E. T., Watson G. N., A course of Modern Analysis, Second edition, 1912.
  • [Wor2000] Woronowicz S.L. Quantum exponential function, Rev. Math. Phys. 12 (2000), 873-920.
  • [Yush1968] Yushkevich A. P. History of mathematics in Russia up to 1917. (Russian) Nauka, Moscow, 1968, 591 pp.
University of Graz;
High School of Modern Mathematics MIPT;
Moscow State University, MechMath Faculty;
Univesity of Vienna, Faculty of Mathematics.
[…]